jack jill and bob own a store together. Jack owns one third, jill owns one quarter what fraction does bob own

Answer by Tutor77(13)

First, you add the fraction that jack and Jill own.
1/3+1/4= 4/12+3/12=7/12
Now, subtract the whole by 7/12.
1-7/12.
Now, get the common denominator, which is 12, and multiply 1/1 by 12.
12/12-7/12=5/12